# A Comprehensive Benchmark for RNA 3D Structure-Function Modeling

### v2

@@ Abstract
The relationship between RNA structure and function has recently attracted interest within the deep learning community, a trend expected to intensify as nucleic acid structure models advance. Despite this momentum, a lack of standardized, accessible benchmarks for applying deep learning to RNA 3D structures hinders progress. To this end, we introduce a collection of seven benchmarking datasets specifically designed to support RNA structure–function prediction. Built on top of the established Python package rnaglib, our library streamlines data distribution and encoding, provides tools for dataset splitting and evaluation, and offers a comprehensive, user-friendly environment for model comparison. The modular and reproducible design of our datasets encourages community contributions and enables rapid customization. To demonstrate the utility of our benchmarks, we report baseline results for all tasks using a relational graph neural network.
Data and code are available at github.com/cgoliver/rnaglib.
@@ 1 Introduction
Recent years have witnessed the advent of deep learning methods for structural biology culminating in the award of the Nobel Prize in Chemistry in 2024. AlphaFold (Jumper et al., 2021) revolutionized protein structure prediction, equipping the field with millions of new structures. Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) or fundamental biology (van Kempen et al., 2022). A key factor in these breakthroughs was the development of neural encoders that directly model protein structure (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022). These advances in turn are rooted in solid competitions (CASP, CAPRI), and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023). By setting clear goals, such benchmarks are the foundation for the development of structure encoders and are required to leverage ever increasing structural data availability. Yet, to date, structure-function benchmarks have focused on proteins, hampering model development for other biological macromolecules.
Ribonucleic acids (RNAs) are a large family among these macromolecules, supporting biological functions along every branch in the tree of life. Besides messenger RNAs which code for proteins, non-coding RNAs are a vast class of molecules which can adopt complex 3D folds (Cech and Steitz, 2014) and take up diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021). However, our understanding of non-coding RNAs and their functions remains limited. This can be largely attributed to the negatively charged nature of RNA backbones, which makes it flexible and limits the availability of high-resolution RNA structures, and imposes significant modeling challenges. Another predominant challenge to a functional understanding of RNA 3D structure lies in the lack of infrastructure for the development and evaluation of function prediction models. In this work, we propose a benchmarking suite to act as a facilitating library for the development of structure encoders on RNA.
Our key contributions include:
@@
  • We apply the benchmark to simple architectures to initialize a leaderboard and investigate the effect of splitting and structural context on performance.
@@ 2.1 Protein Focused Benchmarking
Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005). To our knowledge, ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins. Similar, more comprehensive tools were then proposed, such as ProteinShake (Kucera et al., 2023) ProteinWorkshop (Jamasb et al., 2024) and TorchDrug (Zhu et al., 2022), that unify the above tasks and lower the barrier to develop protein structure encoders. ProteinGym (Notin et al., 2023) addresses the evaluation of protein mutation effects, while Buttenschoen et al. (2024); Kovtun et al. (2024); Durairaj et al. (2024) address protein interactions. These works contain notable efforts to scale datasets using predicted structures and to propose biologically appropriate splitting strategies.
@@ 2.2 RNA Structural Datasets and Benchmarking
In the realm of RNA 3D structure based modeling infrastructure, three papers propose datasets with the objective of facilitating machine learning. They either provide cleaned files in various formats through a web interface (RNAsolo (Adamczyk et al., 2022)), join RNA structures with their corresponding sequence alignment (RNANet (Becquey et al., 2021)) or focus on structure prediction (RNA3DB (Szikszai et al., 2024)). Recently, multiple benchmarking suites were introduced for models on RNA sequences (BEACON (Ren et al., 2024)), predicted structures of short RNAs (Xu et al., 2024), or for fitness prediction (Arora et al., 2025). None of these methods propose benchmark tasks to compare RNA modeling and learning methods on experimental structures.
@@ 2.3 Structure-Based RNA Models
While most deep learning models on RNA focus on secondary structure prediction and sequence-level tasks, some structure-based models have been developed. RBind (Wang et al., 2018) proposed learning on RNA structures using residue graph representations, followed by others integrating sequence features in the learning (Su et al., 2021; Wang et al., 2023). RNAmigos and RNAmigos2 (Oliver et al., 2020; Carvajal-Patiño et al., 2025) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining. ARES (Townshend et al., 2021b) applied a tensor field neural network (Thomas et al., 2018) to RNA structure, modeling it as an atomic graph. Geometric Vector Perceptron (Jing et al., 2021), an equivariant message passing protein encoder, was adapted to RNA in (Joshi et al., 2024; Tan et al., 2023, 2025; Huang et al., 2024; Wong et al., 2024). EquiRNA (Li et al., 2025) proposed a hierarchical architecture. Lately, HARMONY (Xu et al., 2025) introduces a model using multiple RNA modalities.
@@ 3 Tools to assemble tasks
We describe the full process of building a task: data collection, quality filtering, cleanup, and splitting. We developed these tools to build our proposed tasks. They are ready to use and documented for any practitioner wishing to build his own reproducible task with minimal effort.
Data collection and annotation
@@
Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA. In some tasks, filter down the set of RNA-containing PDB structures with the subset proposed in (Leontis and Zirbel, 2012) and later referred to as bgsu. We annotate each system with RNA-level features such as its resolution, and residue-level features such as their coordinates, the presence of interacting compounds, and the amount of protein atoms in the residues’ vicinities.
Structure partitioning and quality filters
@@
RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see figures 5(a) and 5(b)). Many systems have less than 300 residues, while some (mostly rRNAs) have several thousands. We consider independent RNA fragments from the same PDB file should be treated as multiple systems ; we split each of our systems into connected components in the graph formed by RNA residues and their interactions. Concretely, the resulting decrease in system size leads to a reduction in required computing power for downstream steps.
@@
We implement resolution and size filters, with default cutoff values including systems below 4Å resolution and with a length of 15 to 500 residues. In tasks, the upper size cutoff is used to limit computational expense. We also provide a protein content filter which removes RNA structures that are heavily structured through an interaction with a protein. This crucial filter has been overlooked in most existing RNA structural datasets. In addition, we implement the drug-like filter introduced in Hariboss (Panei et al., 2022) for small molecules binding to RNA.
Redundancy filtering and dataset splitting
Biological data points are often related in terms of evolutionary history and topological characteristics. Moreover, as structures of a given system are often observed in slightly different conditions, it is common to obtain copies with minimal variations, introducing redundancy in the data. Insufficient rigor in splitting can thus lead to severe data leaks (Bernett et al., 2024; Volkov et al., 2022). Efforts like the PINDER database (Kovtun et al., 2024) have set a standard in the field for model generalization assessment.
@@
Given a set of RNA molecules $\mathcal{R}$, a similarity metric ${s}$ and a threshold $\theta$, denote as $\mathcal{G}$ the graph $(\mathcal{R},\mathcal{E}^{\theta})$, where $\mathcal{E}^{\theta}:=\{(r_{1},r_{2})\in\mathcal{R}^{2},s(r_{1},r_{2})>\theta\}$. We propose a clustering algorithm that returns connected components of $\mathcal{G}$ as clusters. This ensures that any pair of points in different clusters has a maximal similarity of $\theta$. In our work, we compute sequence-based with CD-HIT (Fu et al., 2012) or structure-based similarity with US-align (Zhang et al., 2022a).
We propose a redundancy removal algorithm that selects the highest resolution system for each cluster. For most tasks, we first apply this algorithm at a sequence similarity threshold of 0.9, and then at a structure similarity threshold of 0.8.
@@
rnaglib includes a data splitting algorithm that aggregates clusters together. To prevent data leakage, we build clusters with a more conservative structural similarity threshold (0.5 unless mentioned otherwise). Then, we cast the aggregation of clusters as an integer programming problem, solved using the linear programming tool PuLP (Mitchell et al., 2011). This enforces structural dissimilarity between splits for predetermined split size ratios while achieving optimal label stratification.
Metrics computation
@@
Given the type of prediction target (e.g. binary classification, multi-label, etc.), each task implements a standardized way to compute evaluation metrics. Each task is evaluated by a recommended metric, specified in Table 2, and a range of additional standard metrics.
Anatomy of a task
@@
A “task” is now simply the collection of the aforementioned components with the choices relevant to a particular biological problem. That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, (ii) redundancy removing and fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol. This modular design will help other practitioners propose additional benchmark tasks, and lower the barrier to entry for training models on RNA structure.
@@ 4 Tasks on RNA 3D Structure
Our tasks suite explores various dimensions of RNA structural research briefly summarized and illustrated in Figure 1. We offer seven tasks among which three are based on previous research with published datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)). For RNA-IF and RNA-Site, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
Armed with an RNA structure, the main focus of each of the subsequent tasks is to either predict aspects of the RNA’s biological function, or to reverse the direction and predict the sequence from which a given structure arises (RNA design). We provide an overview of these tasks from a machine learning point of view relating to the type of task and dataset size in Table 1.
@@
7. RNA-VS Predict the affinity of a small molecule for an RNA binding site, and use it for Virtual Screening (VS).
Figure 1: A graphical overview of the seven tasks included in rnaglib illustrated through a precursor tRNA (4GCW) and a SAM-I riboswitch (5FJC). The integrated table contains a short summary of each task.
@@ 4.1 RNA-GO: Function Tagging
(Definition): This is an RNA-level, multi-label classification task where an RNA is mapped to five possible labels representing molecular functions.
(Context): Effective function prediction models have the capacity to uncover new structure-function connections. The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function to a gene, and thus, indirectly to the RNA or protein it encodes. It resulted in discrete functional categories called GO terms. Predicting the GO term from a protein structure was proposed as a task in (Gligorijević et al., 2021) and has since been regularly used as a benchmark task. A manual annotation of RNA sequences with GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families.
(Processing): We extracted all GO terms corresponding to an RNA structure. Then, we filtered these terms based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), as well as infrequent ones (less than 50 occurrences). We then grouped together GO-terms that were fully correlated, which resulted in five classes. Finally, we discarded RNA fragments with less than 15 residues, resulting in 501 systems. We split these systems based on a sequence similarity cutoff, following Gligorijević et al. (2021).
@@ 4.2 RNA-IF: Molecular Design
(Context): One avenue for designing new macromolecules is to design a coarse grained tertiary structure first, for instance scaffolding a structural motif, and then design a sequence that will fold into this given structure. This second step is denoted as Inverse Folding (IF), since it maps a known structure to an unknown sequence. Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023), with adaptations for RNA with classical models (Leman et al., 2020a). Learning based approaches were pioneered by RDesign (Tan et al., 2023), gRNAde (Joshi et al., 2024) and other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori and Jin, 2024).
(Processing): We gather all connected components in our 3D structure database having sizes between 15 and 300. Then, we remove identical sequences using a redundancy removal step with CD-HIT with a cutoff of 0.99. We cluster the data on structural similarity with a threshold of 0.9, and split these while ensuring a maximal similarity of 0.5. We also provide datasets and splits from gRNAde (Joshi et al., 2024). Our dataset differs by allowing multi-chain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size.

### v3

# A Comprehensive Benchmark for RNA 3D Structure–Function Modeling

@@ Abstract
The relationship between RNA structure and function has recently attracted interest within the deep learning community, a trend expected to intensify as nucleic acid structure models advance. Despite this momentum, the lack of standardized, accessible benchmarks for applying deep learning to RNA 3D structures hinders progress. To this end, we introduce a collection of seven benchmarking datasets specifically designed to support RNA structure–function prediction. Built on top of the established Python package rnaglib, our library streamlines data distribution and encoding, provides tools for dataset splitting and evaluation, and offers a comprehensive, user-friendly environment for model comparison. The modular and reproducible design of our datasets encourages community contributions and enables rapid customization. To demonstrate the utility of our benchmarks, we report baseline results for all tasks using a relational graph neural network.
All code and data, along with documentation are available at rnaglib.org
@@ 1 Introduction
Recent years have witnessed the advent of deep learning methods for structural biology, culminating in the award of the Nobel Prize in Chemistry in 2024. AlphaFold (Jumper et al., 2021) revolutionized protein structure prediction, equipping the field with millions of new structures. Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) and fundamental biology (van Kempen et al., 2022). A key factor in these breakthroughs was the development of neural encoders that directly model protein structure (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022). These advances in turn are rooted in robust competitions (CASP, CAPRI) and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023). By setting clear goals, such benchmarks are the foundation for the development of structure encoders and are required to leverage ever-increasing structural data availability. Yet, to date, structure–function benchmarks have focused on proteins, hampering model development for other equally important biological macromolecules.
Ribonucleic acids (RNAs) are a large family among these macromolecules, supporting biological functions along every branch in the tree of life. Besides messenger RNAs which code for proteins, non-coding RNAs are a vast class of molecules which can adopt complex 3D folds (Cech and Steitz, 2014) and play diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021). However, our understanding of non-coding RNAs and their functions remains limited. This can be largely attributed to the negatively charged nature of RNA backbones, which makes them flexible and limits the availability of high-resolution RNA structures, and poses significant modeling challenges. Another predominant challenge to a functional understanding of RNA 3D structure lies in the lack of infrastructure for the development and evaluation of function prediction models. In this work, we propose a benchmarking suite to serve as a facilitating library for the development of structure encoders on RNA.
Our key contributions are:
@@
  • We apply the benchmark to simple architectures to initialize a leaderboard and investigate the effect of splitting, structural context, and RNA representation choice on performance.
@@ 2.1 Protein Focused Benchmarking
Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005). ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins. More comprehensive tools were proposed, such as ProteinShake (Kucera et al., 2023) ProteinWorkshop (Jamasb et al., 2024) and TorchDrug (Zhu et al., 2022), that unify the above tasks and lower the barrier to develop protein structure encoders. ProteinGym (Notin et al., 2023) addresses the evaluation of protein mutation effects, while Buttenschoen et al. (2024); Kovtun et al. (2024); Durairaj et al. (2024) address protein interactions. These works are notable efforts to scale datasets using predicted structures and to propose biologically appropriate splitting strategies.
@@ 2.2 RNA Structural Datasets and Benchmarking
In RNA 3D structure-based modeling infrastructure, three papers propose machine learning ready datasets. They either provide cleaned files in various formats through a web interface (RNAsolo (Adamczyk et al., 2022)), join RNA structures with their sequence alignment (RNANet (Becquey et al., 2021)) or focus on structure prediction (RNA3DB (Szikszai et al., 2024)). Recently, multiple benchmarking suites were introduced for models on RNA sequences (BEACON (Ren et al., 2024)), predicted structures of short RNAs (Xu et al., 2024), or for fitness prediction (Arora et al., 2025). None of these methods propose benchmark tasks to compare RNA modeling and learning methods on experimental structures. Our effort is integrated within a previously published python package, rnaglib (Mallet et al., 2022) that facilitates the use of RNA structural data with 2.5D graphs. This work adds all necessary material to process data and datasets, as well as a set of seven tasks (see Supplementary Section A for a more detailed comparison).
@@ 2.3 Structure-Based RNA Models
Whilst most deep learning models on RNA focus on secondary structure prediction and sequence-level tasks, some structure-based models were developed. RBind (Wang et al., 2018) proposed learning on RNA structures using residue graph representations, followed by others integrating sequence features in the learning (Su et al., 2021; Wang et al., 2023). RNAmigos (Oliver et al., 2020) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining. ARES (Townshend et al., 2021b) applied a tensor field neural network (Thomas et al., 2018) to RNA structure, modeling it as an atomic graph. Geometric Vector Perceptron (Jing et al., 2021), an equivariant message passing protein encoder, was adapted to RNA in (Joshi et al., 2024; Tan et al., 2023; 2025; Huang et al., 2024; Wong et al., 2024). EquiRNA (Li et al., 2025) proposed a hierarchical architecture. Recently, HARMONY (Xu et al., 2025) introduces a model using multiple RNA modalities.
@@ 3 Tools to assemble tasks
We describe the full process of building a task: data collection, quality filtering, cleanup, and splitting. Tasks are ready to use and documented for any practitioner wishing to build their own reproducible tasks with minimal effort.
@@
Data collection and annotation Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA. In some tasks, we use the non-redundant structures subset proposed in (Leontis and Zirbel, 2012) and later referred to as bgsu. We annotate each system with RNA-level features such as its resolution, and residue-level features such as coordinates, presence of interacting compounds or number of protein atoms in the residues’ vicinities.
@@
Structure partitioning and quality filters RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see Supplementary Figures 5(a) and 5(b)). Many systems have less than 300 residues, while some (mostly rRNAs) have several thousands. We treat RNA fragments from the same PDB file as multiple systems if they do not interact, decreasing system size and reducing required computing power for downstream steps.
@@
We implement resolution and size filters, with default cutoffs including systems below 4Å resolution and between 15 to 500 residues; the upper size cutoff limits computational expense. A novel protein content filter removes RNA structures heavily dependent on protein interactions, a crucial step overlooked in most existing datasets. We also include the drug-like filter from Hariboss (Panei et al., 2022) to screen small molecules binding to RNA.
@@
Redundancy filtering and dataset splitting Rigorous data splitting is often key for structural biology benchmarks such as the PINDER database (Kovtun et al., 2024) for proteins, which has set a standard in the field for model generalization assessment. Given a set of RNA molecules $\mathcal{R}$, a similarity metric s{s} and a threshold $\theta$, let $\mathcal{G}$ be the graph connecting similar RNAs, $\mathcal{G}=(\mathcal{R},\mathcal{E}^{\theta})$, where $\mathcal{E}^{\theta}:=\{(r_{1},r_{2})\in\mathcal{R}^{2},s(r_{1},r_{2})>\theta\}$. We propose a clustering algorithm that returns connected components of $\mathcal{G}$ as clusters, ensuring that points in different clusters have a maximal similarity of $\theta$. We compute similarity s{s} with CD-HIT (Fu et al., 2012) for sequence and with US-align (Zhang et al., 2022a) for structure.
@@
Building on this, we propose a redundancy removal step by selecting the highest-resolution RNA in each cluster. For most tasks, we use a 0.9 sequence similarity threshold and then 0.8 for structural similarity. We also propose a data splitting algorithm that aggregates clusters. To prevent data leakage, we build dissimilar clusters with a conservative structural similarity threshold (0.5 unless mentioned otherwise). Then, we cast the cluster aggregation into splits of given sizes as an integer programming problem, solved with the linear programming tool PuLP (Mitchell et al., 2011); this enforces structural dissimilarity and allows using label stratification.
@@
Metrics computation Given the type of prediction target (e.g., binary classification, multilabel, etc.), each task implements a standardized way to compute evaluation metrics. Each task is evaluated by a recommended metric, specified in Table 1, and a range of additional standard metrics.
@@
Anatomy of a task A “task” is now simply the collection of the aforementioned components with the choices relevant to a particular biological problem. That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, (ii) redundancy removing and fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol. This modular design will help other practitioners propose additional benchmark tasks, and lower the barrier to entry for training models on RNA structure.
@@ 4 Tasks on RNA 3D Structure
Our tasks suite explores various dimensions of RNA structural research illustrated and briefly summarized in Figure 1. We offer seven tasks among which three are based on previous research with published datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)). For RNA-IF and RNA-Site, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
Armed with an RNA structure, the focus of each of the tasks is to predict aspects of the RNA’s biological function, or to reverse the direction and predict the sequence from which a given structure arises (RNA design). Further information on task datasets is found in Table 1 and supplement C.
@@
7. RNA-VS Predict small molecule-RNA binding site affinity, and use it for Virtual Screening (VS).
Figure 1: A graphical overview of the seven tasks included in our library, illustrated by a precursor tRNA (4GCW) and a SAM-I riboswitch (5FJC). The integrated table contains short task summaries.
@@ 4.1 RNA-GO: Function Tagging
(Definition): This is an RNA-level, multilabel classification task where an RNA is mapped to five possible labels representing molecular functions.
(Context): Function-prediction models have the capacity to uncover new structure-function connections. The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function with a gene, and thereby to the RNA or protein it encodes. It resulted in functional categories called GO terms, whose prediction from a protein structure was proposed as a task in (Gligorijević et al., 2021), which is now regularly used. A manual annotation of RNAs with GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families.
(Processing): GO terms corresponding to an RNA structure are extracted, then filtered based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), and infrequent ones (fewer than 50 occurrences). We group together GO-terms that are fully correlated, resulting in five classes. Finally, we discard RNA fragments with fewer than 15 residues, resulting in 499 systems. We split the data along sequence similarity, following Gligorijević et al. (2021).
@@ 4.2 RNA-IF: Molecular Design
(Context): One avenue for designing new macromolecules is to fix a coarse-grained tertiary structure first, for instance scaffolding a structural motif, and then predict a sequence that will fold into this given structure. This second step is called inverse folding (IF), since it maps a known structure to an unknown sequence. Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023), with adaptations for RNA with classical models (Leman et al., 2020a). Learning-based approaches were pioneered by RDesign (Tan et al., 2023), gRNAde (Joshi et al., 2024), while other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori and Jin, 2024).
(Processing): We gather all connected components in our database between 15 and 300 nucleotides, then remove identical sequences using CD-HIT redundancy removal with a threshold of 0.9 and split on structural similarity at a threshold of 0.5. We also provide datasets and splits from gRNAde (Joshi et al., 2024). Our dataset differs by allowing multichain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size.